To study bonding between mothers and infants, a researcher places each mother and her infant in a playroom and has the mother leave for 10 minutes. The researcher records crying time in the sample of infants during this time that the mother was not present and finds that crying time is normally distributed with

M = 5

and

SD = 1.2.

Based on the empirical rule, state the range of crying times within 68% of infants cried, 95% of infants cried, and 99.7% of infants cried.

(a) 68% of infants cried
to  min

(b) 95% of infants cried
to  min

(c) 99.7% of infants cried
to  min
Answer #1

Given that, mean = M = 5 and standard deviation = SD = 1.2

According to Empirical rule,

i) Approximately 68% of the data fall within 1 standard deviations of the mean.

(mean - sd = 5 - 1.2 = 3.8 and mean + sd = 5 + 1.2 = 6.2)

ii) Approximately 95% of the data fall within 2 standard deviations of the mean.

(mean - 2sd = 5 - (2 * 1.2) = 2.6 and mean + 2sd = 5 + (2 * 1.2) = 7.4)

iii) Approximately 99.7% of the data fall within 3 standard deviations of the mean.

(mean - 3sd = 5 - (3 * 1.2) = 1.4 and mean + 3sd = 5+ (3 * 1.2) = 8.6)

Therefore,

a) 68% of infants cried 3.8 to 6.2 min.

b) 95% of infants cried 2.6 to 7.4 min.

c) 99.7% of infants cried 1.4 to 8.6 min.